# The NetMob23 Dataset: Population Density and OD Matrices from Four LMIC Countries

### v1

@@ Abstract
The NetMob24 dataset offers a unique opportunity for researchers from a range of academic fields to access comprehensive spatiotemporal data sets spanning four countries (India, Mexico, Indonesia, and Colombia) over the course of two years (2019 and 2020). This dataset, developed in collaboration with Cuebiq (Also referred to as Spectus)111https://www.cuebiq.com/, comprises privacy-preserving aggregated data sets derived from mobile application (app) data collected from users who have voluntarily consented to anonymous data collection for research purposes. It is our hope that this reference dataset will foster the production of new research methods and the reproducibility of research outcomes.
@@ 1 Introduction
In light of the widespread adoption of mobile technology, the digital data generated from mobile devices and mobile applications provides us with the ability to examine a multitude of human behaviors, including mobility, at a previously unattainable scale. These data have the potential to facilitate the analysis of the habits of large populations at large scales, such as those of cities or countries. In certain instances, they can even supplant traditional sources of data, such as surveys in countries where data are scarce. The value of data for development and enhancement of public policies is largely untapped, as they can help fill data gaps and provide real-time and finer-scale insight (WB, 2021). While there are clear benefits to be gained from leveraging these data for the public good, there are also a number of challenges to be addressed. These include the need to ensure data privacy and security, as well as the need to develop new methodologies for data analysis. In the context of data for public good, several topics emerge as particularly relevant, including the use of mobile data for transportation planning, disaster recovery and epidemic response, socioeconomic analysis, and tourism. In this paper, we start out by providing an overview of the current state of research in these areas, highlighting the key findings and challenges that remain to be addressed. We then provide an overview of the Data Challenge dataset that has been made available to facilitate new research on these and other topics. We present some of the limitations of this data as well as some descriptive analysis. Finally, we highlight additional data sources that could be combined with the data to help in studying some of these topics.
@@
With population growth in many developing countries and concentration of resources and opportunities in urban areas, many cities around the world are facing challenges in terms of transportation. The use of mobile phone data can provide valuable insights into the movement of populations in urban areas, which can be used to inform the development of transportation infrastructure and services. For example, in González et al. (2008), the authors use mobile phone data to study the movement of people in cities, with the aim of improving the efficiency of public transportation systems. Different mobile phone data is used to study the movement of people in urban area, with the aim of improving the design of transportation networks including GPS traces from mobile phones (Pappalardo et al., 2015), mobile phone traffic data (Furno et al., 2016), mobile network metadata (Khodabandelou et al., 2019), and mobile phone data (Bachir et al., 2019; Vilella et al., 2020). Finally, Ucar et al. (2021) use mobile phone metadata to study the relationship between the use of mobile services and the development of transportation infrastructure in cities.
@@
Disaster recovery represents a critical domain where targeted resource allocation and response during crises is of paramount importance. The provision of a dynamic view of the situation, with the objective of enhancing situational awareness, represents a key area for consideration, as it allows for a more targeted and effective response. In this context, mobile phone data of all kinds can provide valuable insights into the movement of populations during and after disasters, as well as the impact of these movements Yabe et al. (2021); Lu et al. (2012); Wang and Taylor (2014); Bagrow et al. (2011), for example, on the demand for health services or the pre-positioning of response operations Wang et al. (2021). In addition to mobile data, other proxies for human mobility can also provide valuable information. For instance, Alatrista-Salas et al. (2021) and Eyre et al. (2020) analyze the impact of natural disasters on purchasing behavior, with a particular focus on shifts in healthcare demand to alternative facilities. For a more detailed examination of the role played by mobile phone location data in the disaster recovery process, please refer to Yabe et al. (2022).
@@
In recent studies of the tourism industry, various methodologies have been proposed to enhance the understanding of tourist behaviors and flows. Kovalcsik et al. (2022) introduced a methodology aimed at comprehending tourist flows by accounting for unobserved tourists. Similarly, Altin et al. (2022) utilized Call Detail Record (CDR) data to analyze the different types of visits to Estonia from 2006 to 2013. Expanding on the use of CDR data, the work of Xu et al. (2021) compares tourist mobility patterns across various cities in South Korea. Grassini et al. (2021) focused on analyzing the volume of tourist flows to Florence, Italy. Park et al. (2023) proposed a model for segmenting tourists based on their activities using mobile phone data. In Hungary, Michalkó et al. (2023) examined tourism dynamics between large cities and their surrounding areas using mobile phone data. Lastly, Sun et al. (2021) developed a methodology to differentiate tourists from locals within extensive mobile phone data sets. Together, these studies illustrate a growing trend in leveraging mobile and CDR data to refine our insights into tourist behaviors and patterns.
@@
This article is organized as follows. Section 2 reviews the data sources used to generate the netmob24 dataset. Section 3 presents the aggregation methodology and ethical considerations used for the generation of this dataset. Section 4 we provide a details description of the datasets provided, and the possible anomalies present inside it. Finally, we conclude we some additional resources in section 5.
@@ 2.1 Device Location Data
The device location data records the location of a given device at specific times. This dataset includes details such as the event time, anonymous ID, coordinates, accuracy in meters, operating system name, device manufacturer name, timezone offset in seconds, and speed in meters per second, among other variables. It should be noted that in some cases the original latitude and longitude values are transformed to preserve the privacy of users. For example, home areas are re-assigned to centroid of the corresponding Geohash 6 tile, and points falling with Sensitive Points of Interest are removed from the dataset for privacy purposes. Cuebiq first classifies the point as a recurring area, a whitelisted area222These are points of interest (POI) included in Cuebiq’ POI whitelist such as commercial locations and other POIs that are permissible for use according to Cuebiq’ privacy requirements. or an ”Other” area. Any point that is in a recurring area is transformed to the coordinates of the centroid of the nearest geometry with 600+ households. Importantly, the point always remains in the same geohash 6 where the original point was located. For generating the population dataset, we use only the event time, anonymous ID, and coordinates variables. Data points with latitude or longitude values recorded as errors or zeros were removed.
@@ 4 Data Description
This section presents an exploratory analysis of the provided data challenge PD and OD datasets. The analysis aims to offer an understanding of the datasets from the points of view of completeness, spatial, and time interval perspectives, highlighting their strengths and weaknesses, to support participants with their research applications.
@@ 4.1 Data Completeness and Anomalies
Figure 4 presents the time series of the number of unique users from the PD dataset, the trip count from the OD dataset, and the ratio of trip count to unique users for 2019 and 2020. This visualization aims to highlight key feature changes and identify anomalies. Notably, visually, some of the days we can observe anomalies include May 10 to May 20, 2019, as well as on October 22, 2019. We are identifying these anomalies visually based on drops in the line charts in Figure 4(a) and Figure 4(b), but when conducting analysis with the data, it may be advisable to identify all anomalies using a standard definition. An example of such a definition would be identifying observations that are more than 2 standard deviations from the average values within 30 days. It is important to note that for the OD data, certain days are missing data for certain countries. The OD has missing data for all four countries on 20200501; India and Indonesia for 20191022 and 20191231. Additionally, there are some days where certain time intervals are missing (e.g. Colombia and Mexico on 20191231 for the hourly intervals starting at 0,3,6, and 9). In the 3h datasets, these observations would be missing, but in the daily datasets, there would be an observation for the day since there are some time intervals with data, but the values would be much lower since part of the day is missing. Again, these cases could be identified through anomalie detection methods.
@@
So far the focus has been on aspects of the data generating process that may impact the data at a particular time, but another important area related to completeness is the representativeness of the data. In order to be part of the sample of users in the data, it is a necessary condition to have a smartphone and to also have the resources to pay for mobile data to use the apps through which the data points are generated. When studying low and middle income countries, there can be large portions of the population that are not able to afford a smartphone or to pay for mobile data on a regular basis, and therefore would not be represented in the dataset Milusheva et al. (2021a). Additionally, the apps that different people use may differ based on demographic characteristics like gender or income, and as some apps become excluded or included in contributing to the data, the demographic make-up of those contributing their data could change, affecting also the behaviors seen in the data. This can also vary from country to country as certain apps may be more popular in some countries and not in others. These are not aspects that it is possible to correct for in the datasets provided for the Challenge, but important aspects to consider when interpreting the data and results and to be further analyzed in the future.

### v2

# The NetMob2024 Dataset: Population Density and OD Matrices from Four LMIC Countries

@@ Abstract
The NetMob24 dataset offers a unique opportunity for researchers from a range of academic fields to access comprehensive spatiotemporal data sets spanning four countries (Colombia, Indonesia, India, and Mexico) over the course of two years (2019 and 2020). This dataset, developed in collaboration with Cuebiq (Also referred to as Spectus)111https://www.cuebiq.com/, comprises privacy-preserving aggregated data sets derived from mobile application (app) data collected from users who have voluntarily consented to anonymous data collection for research purposes. It is our hope that this reference dataset will foster the production of new research methods and the reproducibility of research outcomes.
@@ 1 Introduction
In light of the widespread adoption of mobile technology, the digital data generated from mobile devices and mobile applications provides us with the ability to examine a multitude of human behaviors, including mobility, at a previously unattainable scale. These data have the potential to facilitate the analysis of the habits of large populations at large scales, such as those of cities or countries. In certain instances, they can even supplant traditional sources of data, such as surveys in countries where data are scarce. The value of data for development and enhancement of public policies is largely untapped, as they can help fill data gaps and provide real-time and finer-scale insights (WB, 2021). While there are clear benefits to be gained from leveraging these data for the public good, there are also a number of challenges to be addressed. These include the need to ensure data privacy and security, as well as the need to develop new methodologies for data analysis. In the context of data for public good, several topics emerge as particularly relevant, including the use of mobile data for transportation planning, disaster recovery and epidemic response, socioeconomic analysis, and tourism. In this paper, we start out by providing an overview of the current state of research in these areas, highlighting the key findings and challenges that remain to be addressed. We then provide an overview of the Data Challenge dataset that has been made available to facilitate new research on these and other topics. We present some of the limitations of this data as well as some descriptive analysis. Finally, we highlight additional data sources that could be combined with the data to help in studying some of these topics.
@@
With population growth in many developing countries and concentration of resources and opportunities in urban areas, many cities around the world are facing challenges in terms of transportation. The use of mobile phone data can provide valuable insights into the movement of populations in urban areas, which can be used to inform the development of transportation infrastructure and services. For example, in González et al. (2008), the authors use mobile phone data to study the movement of people in cities, with the aim of improving the efficiency of public transportation systems. Different mobile phone data is used to study the movement of people in urban areas with the aim of improving the design of transportation networks including GPS traces from mobile phones (Pappalardo et al., 2015), mobile phone traffic data (Furno et al., 2016), mobile network metadata (Khodabandelou et al., 2019), and mobile phone data (Bachir et al., 2019; Vilella et al., 2020). Finally, Ucar et al. (2021) use mobile phone metadata to study the relationship between the use of mobile services and the development of transportation infrastructure in cities.
@@
Disaster recovery represents a critical domain where targeted resource allocation and response during crises is of paramount importance. The provision of a dynamic view of the situation, with the objective of enhancing situational awareness, represents a key area for consideration, as it allows for a more targeted and effective response. In this context, mobile phone data of all kinds can provide valuable insights into the movement of populations during and after disasters, as well as the impact of these movements Yabe et al. (2021); Lu et al. (2012); Wang and Taylor (2014); Bagrow et al. (2011), for example, on the demand for health services or the pre-positioning of response operations (Wang et al., 2021). In addition to mobile data, other proxies for human mobility can also provide valuable information. For instance, Alatrista-Salas et al. (2021) and Eyre et al. (2020) analyze the impact of natural disasters on purchasing behavior, with a particular focus on shifts in healthcare demand to alternative facilities. For a more detailed examination of the role played by mobile phone location data in the disaster recovery process, please refer to Yabe et al. (2022).
@@
In recent studies of the tourism industry, various methodologies have been proposed to enhance the understanding of tourist behaviors and flows. Kovalcsik et al. (2022) introduced a methodology aimed at comprehending tourist flows by accounting for unobserved tourists. Similarly, Altin et al. (2022) utilized Call Detail Record (CDR) data to analyze the different types of visits to Estonia from 2006 to 2013. Expanding on the use of CDR data, the work of Xu et al. (2021) compares tourist mobility patterns across various cities in South Korea. Grassini et al. (2021) focus on analyzing the volume of tourist flows to Florence, Italy. Park et al. (2023) proposed a model for segmenting tourists based on their activities using mobile phone data. In Hungary, Michalkó et al. (2023) examined tourism dynamics between large cities and their surrounding areas using mobile phone data. Lastly, Sun et al. (2021) developed a methodology to differentiate tourists from locals within extensive mobile phone data sets. Together, these studies illustrate a growing trend in leveraging mobile and CDR data to refine our insights into tourist behaviors and patterns.
@@
This article is organized as follows. Section 2 reviews the data sources used to generate the netmob24 dataset. Section 3 presents the aggregation methodology and ethical considerations used for the generation of this dataset. Section 4 provides a detailed description of the datasets, and the possible anomalies present within the data. Finally, we conclude with some additional resources in section 5.
@@ 2.1 Device Location Data
The device location data records the location of a given device at specific times. This dataset includes details such as the event time, anonymous ID, coordinates, accuracy in meters, operating system name, device manufacturer name, timezone offset in seconds, and speed in meters per second, among other variables. It should be noted that in some cases, the original latitude and longitude values are transformed to preserve the privacy of users. For example, home areas are re-assigned to the centroid of the corresponding Geohash 6 tile, and points falling under Sensitive Points of Interest are removed from the dataset for privacy purposes. Cuebiq first classifies the point as a recurring area, a whitelisted area222These are points of interest (POI) included in Cuebiq’ POI whitelist such as commercial locations and other POIs that are permissible for use according to Cuebiq’ privacy requirements. or an ”Other” area. Any point that is in a recurring area is transformed to the coordinates of the centroid of the nearest geometry with 600+ households. Importantly, the point always remains in the same geohash 6 where the original point was located. For generating the population dataset, we use only the event time, anonymous ID, and coordinates variables. Data points with latitude or longitude values recorded as errors or zeros were removed.
@@ 4 Data Description
This section presents an exploratory analysis of the provided data challenge PD and OD datasets. The analysis aims to offer an understanding of the datasets from the point of view of completeness, spatial, and time interval perspectives, highlighting their strengths and weaknesses, to support participants with their research applications.
@@ 4.1 Data Completeness and Anomalies
Figure 4 presents the time series of the number of unique users from the PD dataset, the trip count from the OD dataset, and the ratio of trip count to unique users for 2019 and 2020. This visualization aims to highlight key feature changes and identify anomalies. Visually, some of the days we can observe anomalies include May 10 to May 20, 2019, as well as on October 22, 2019. We are identifying these anomalies visually based on drops in the line charts in Figure 4(a) and Figure 4(b), but when conducting analysis with the data, it may be advisable to identify all anomalies using a standard definition. An example of such a definition would be identifying observations that are more than 2 standard deviations from the average values within 30 days. It is important to note that for the OD data, certain days are missing data for certain countries. The OD has missing data for all four countries on 20200501; India and Indonesia for 20191022 and 20191231. Additionally, there are some days where certain time intervals are missing (e.g. Colombia and Mexico on 20191231 for the hourly intervals starting at 0,3,6, and 9). In the 3h datasets, these observations would be missing, but in the daily datasets, there would be an observation for the day since there are some time intervals with data, but the values would be much lower since part of the day is missing. Again, these cases could be identified through anomalie detection methods.
@@
So far the focus has been on aspects of the data generating process that may impact the data at a particular time, but another important area related to completeness is the representativeness of the data. In order to be part of the sample of users in the data, it is a necessary condition to have a smartphone and to also have the resources to pay for mobile data to use the apps through which the data points are generated. When studying low- and middle-income countries, there can be large parts of the population that are not able to afford a smartphone or to pay for mobile data on a regular basis, and therefore would not be represented in the dataset (Milusheva et al., 2021a). Additionally, the apps that different people use may differ based on demographic characteristics like gender or income, and as some apps become excluded or included in contributing to the data, the demographic make-up of those contributing their data could change, affecting also the behaviors seen in the data. This can also vary from country to country as certain apps may be more popular in some countries and not in others. These are not aspects that it is possible to correct for in the datasets provided for the Challenge, but important aspects to consider when interpreting the data and results and to be further analyzed in the future.